40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 14 10:00:14 2010, josephb, Update, Phase Camera, Work near 1Y2 yesterday 
    Reply  Wed Jul 14 11:51:48 2010, Razib, Update, Phase Camera, Work near 1Y2 yesterday 
Message ID: 3215     Entry time: Wed Jul 14 11:51:48 2010     In reply to: 3213
Author: Razib 
Type: Update 
Category: Phase Camera 
Subject: Work near 1Y2 yesterday 

Quote:

Razib and I were attempting to get the output of a photodiode (PD55A in this case) recorded, so that we could independently measure the slow (~1-10 Hz) fluctuations of the light incident on the camera.  This would then allow us to subtract those fluctuations out, letting us get at the camera noise in the case with signal present (as opposed to just a dark noise measurement when we look at the noise with no signal present).

Originally I was thinking of using one empty patch panel BNCs used for PEM channels down by the 1Y7 rack and go through a 110B, although Alberto pointed out he had recently removed some monitoring equipment, which watched the amplitude modulation at various frequencies of the RF distribution (i.e. 33 MHz, etc).  This equipment output a DC voltage proportional to the amplitude of the RF signals.  The associated channel names were C1:IOO-RFAMPD_33MHZ, C1:IOO-RFAMPD_33MHZ_CAL, C1:IOO-RFAMPD_133MHZ, etc.  These are slow channels, so I presume they enter in via the slow computers, probably via pentek (I didn't check that, although in hindsight I probably should have taken the time to find exactly where they enter the system).  The connections them selves were a set of BNCs on the south side, half way up the 1Y2 rack.

We simply chose one, the 33 MHz channel in this case, and connected.  At around this time, the MC did become unlocked, although it looked like it was due to the MC2 watchdog tripping.  The initial theory was we had bumped the Mode Cleaner while looking around for some BNC cables, although from what Rana had to do last night, it probably was the connection.  We were able to restore the watchdog and confirm that the optic started to settle down again.  Unfortunately, I had to leave about 5 minutes later, and didn't do as thorough an investigation as was warranted.

 Before I left, I disconnected the PD55, so the 33 MHz channel wasn't physically connected to anything!!! Only one end of the wire was connected to the rack while the other was free...

So it wasn't the PD connection that is responsible for MC tripping at the later time...

ELOG V3.1.3-